Calculate the % sodium in a 0.721 g detergent sample. 

Final yield is 0.320 g NaP2O4

1
Expert's answer
2021-02-09T03:46:48-0500

Moles of NaP2O4 = 0.320/149

= 0.00215 mol

So, moles of Sodium is also = 0.00215 mol in NaP2O4

As we know according to law of conservation of mass,

Moles of sodium in detergent is also = 0.00215 mol

So, mass of Sodium = mole × molar mass

= 0.00215 × 23 = 0.0494 g

%Sodium = "Mass of Sodium\u00d7100\/Total Mass"

= 0.0494×100/0.721

= 6.85 %

% Sodium = [6.85] %